OpenStudy (anonymous):

6. Is the system consistent or inconsistent? y = -1/2x - 1 and y = 3x + 3 (Points : 2) Consistent Inconsistent Cannot be determined

OpenStudy (bahrom7893):

can u define consistent/inconsistent? first time im hearing the terms..

OpenStudy (anonymous):

If you say a system is consistent, it means that it can be solved. If you say a system is inconsistent, it means that the system has no solution.

OpenStudy (bahrom7893):

ohh i see.. why couldn't they say so lol.. okay let's check

OpenStudy (bahrom7893):

just had to post that.. back to the problem

OpenStudy (anonymous):

lol

OpenStudy (bahrom7893):

y = -1/2x - 1 and y = 3x + 3 set them equal to each other and try solving: (-1/2)x - 1 = 3x+3 x + 2 = -6x - 6 x+6x = -6-2 7x = -8 x = -8/7 I can solve it, so it must be consistent.

OpenStudy (anonymous):

thanks!